4. IfP(A) 0.3, P(B) 0.2, and P(An B) 0., determine the following probabilities: (b) P(AU B) (c) P(An B) (d) P(An B) (e) P[(A U B)] (f) P(AU B) 5. The following table summarizes the analysis of samples of galvanized steel for coating weight and surface roughness: Coating Weight High High Low Low 16 34 Surface 12 Roughness If the coating weight of a sample is high, what is the probability that the surface roughness is high? If the surface roughness of a sample is high, what is the probability that the coating weight is high? If the surface roughness of a sample is low, what is the probability that the coating weight is low? What is your space of events? a. b. c. d.
Answer #1

4.

P(A)=0.3 , P(B)=0.2 ,  P(A n B) = 0.1

a)

P(A')=1-P(A) =1-0.3

=0.7

b)

P(AU B) = P(A) + P(B) _ P(An B)

=0.3+0.2-0.1

=0.4

c)

P(A'cap B)=P(B)-P(Acap B)=.2-.1

=.1

d)

P(Acap B')=P(A)-P(Acap B)=0.3-0.1

=0.2

e)

P[(Acup B)']=1-P(Acup B)

=1-0.4

=0.6

f)

P(A'cup B)=P(A')+P(B)-P(A'cap B)

=0.7+0.2-0.1

=0.8